# No Metric to Rule Them All: Toward Principled Evaluations of Graph-Learning Datasets

### v1

@@ Abstract
Benchmark datasets have proved pivotal to the success of graph learning, and good benchmark datasets are crucial to guide the development of the field. Recent research has highlighted problems with graph-learning datasets and benchmarking practices—revealing, for example, that methods which ignore the graph structure can outperform graph-based approaches on popular benchmark datasets. Such findings raise two questions: (1) What makes a good graph-learning dataset, and (2) how can we evaluate dataset quality in graph learning? Our work addresses these questions. As the classic evaluation setup uses datasets to evaluate models, it does not apply to dataset evaluation. Hence, we start from first principles. Observing that graph-learning datasets uniquely combine two modes—the graph structure and the node features—, we introduce Rings, a flexible and extensible mode-perturbation framework to assess the quality of graph-learning datasets based on dataset ablations—i.e., by quantifying differences between the original dataset and its perturbed representations. Within this framework, we propose two measures—performance separability and mode complementarity—as evaluation tools, each assessing, from a distinct angle, the capacity of a graph dataset to benchmark the power and efficacy of graph-learning methods. We demonstrate the utility of our framework for graph-learning dataset evaluation in an extensive set of experiments and derive actionable recommendations for improving the evaluation of graph-learning methods. Our work opens new research directions in data-centric graph learning, and it constitutes a first step toward the systematic evaluation of evaluations.
@@ 1 Introduction
    Experiments. We demonstrate the value of our framework and measures through extensive experiments on real-world graph-classification datasets.
@@ 2.1 Mode Perturbations
Given a dataset $\operatorname{\mathcal{D}}$ of attributed graphs, a dataset perturbation is given by $\operatorname{\Phi}(\operatorname{\mathcal{D}}):=\{\operatorname{\varphi}(G,X):(G,X)\in\operatorname{\mathcal{D}}\}$.
@@ 2.2 Performance Separability
Intuitively, given two perturbations $\operatorname{\Phi}$, $\operatorname{\Phi}^{\prime}$ of a dataset $\operatorname{\mathcal{D}}$ as well as a task to be solved on $\operatorname{\mathcal{D}}$, performance separability measures the distance between performance distributions associated with $\operatorname{\Phi}(\operatorname{\mathcal{D}})$ and $\operatorname{\Phi}^{\prime}(\operatorname{\mathcal{D}})$. For a formal definition, we need notation describing these distributions.
@@
To elucidate the relationship between the graph structure and the node features as it pertains to performance, within Rings, we tune models not only on datasets $\operatorname{\mathcal{D}}$ but also on perturbed datasets $\operatorname{\Phi}(\operatorname{\mathcal{D}})$.
@@
For a mode perturbation $\operatorname{\varphi}$, $\operatorname{\operatorname{\mathcal{M}}_{\operatorname{\varphi}}}\coloneqq(\operatorname{\Phi}(\operatorname{\mathcal{D}}),\operatorname{\mathcal{A}},\operatorname{\theta})$ denotes a model tuned to solve the task associated with $\operatorname{\mathcal{D}}$ under mode perturbation $\operatorname{\varphi}$.
@@ 3 Related Work
Figure 3: Comparing GNN performance across different versions of the same dataset. We show the mean (dot) and $95$th percentile interval (bars) of accuracy and AUROC across $100$ runs of the best (as measured by the respective performance mean) among our tuned GAT, GCN, and GIN models, for the original version and $5$ perturbations of $13$ graph-learning datasets. Black resp. silver horizontal lines show best mean trained resp. untrained performance on the original data. For Reddit-M, the complete-graph and complete-features perturbations failed to train due to memory problems, but the existing results already allow us to conclude that this dataset lacks performance separability.
@@ 4 Experiments
We now demonstrate how to use Rings to evaluate the quality of graph-learning datasets, guided by the principles developed in Section 1. Leveraging our mode-perturbation framework, we explore P1 via performance separability and P2 via mode complementarity, before distilling our observations into an actionable taxonomy of recommendations. For further details and additional results, see Appendix D. [^3] We make all code, data, results, and workflows publicly available. Upon publication, a fully documented reproducibility package will be deposited via Zenodo.
Evaluated Datasets. In our main experiments, we evaluate $13$ popular graph-classification datasets: From the life sciences, we select AIDS, ogbg-molhiv (MolHIV), MUTAG, and NCI1 (small molecules), as well as DD, Enzymes, Peptides-func (Peptides), and PROTEINS-full (Proteins) (larger chemical structures). From the social sciences, we take COLLAB, IMDB-B, and IMDB-M (collaboration ego-networks), as well as REDDIT-B and REDDIT-M (online interactions). See Appendix C for details and references.
@@ 4.1 Performance Separability (P1)
For each version of each dataset, we identify the model with the best performance under evaluation metric $\operatorname{f}\in$accuracy, AUROC} as $(\operatorname{\mathcal{A}}^{\star},\operatorname{\theta}^{\star})\coloneqq\arg\max_{(\operatorname{\mathcal{A}},\operatorname{\theta})}\mathbb{E}\left[\operatorname{\operatorname{f}(\operatorname{\operatorname{\mathcal{M}}_{\operatorname{\varphi}}};\operatorname{\zeta})}\right]$, using the performance distribution of the top model, $\operatorname{\mathcal{M}}^{\star}_{\operatorname{\varphi}}\coloneqq(\operatorname{\Phi}(\operatorname{\mathcal{D}}),\operatorname{\mathcal{A}}^{\star},\operatorname{\theta}^{\star})$, to assess performance separability.
@@ 5 Discussion
Conclusions. We introduced Rings, a principled framework for evaluating graph-learning datasets based on mode perturbations—i.e., controlled changes to their graph structure and node features. In Rings, we developed performance separability as a task-dependent, and mode complementarity as a task-independent measure of dataset quality. Using our framework to categorize $13$ popular graph-classification datasets, we identified several benchmarks that require realignment (e.g., by better data modeling) or complete deprecation. While some datasets have been observed to be problematic in prior work, with Rings, we offer a full perturbation-based evaluation pipeline. We designed our initial set of perturbations from a data-centric perspective, but cleverly crafted perturbations could also isolate and confirm the performance separability of any proposed contribution. Thus, we hope that Rings will raise the standard of proof for model-centric evaluation in graph learning as well.
@@
Limitations. Given that message passing is the predominant paradigm in graph learning, our performance-separability experiments target message-passing GNNs, at the exclusion of other architectures. While Rings is a general framework, our current norm-based comparison of metric spaces assumes that nodes and node features are paired. Thus, we presently do not measure aspects like the metric distortion arising from different mappings. Moreover, our notion of mode complementarity is deliberately task-independent and model-agnostic. This allows us to gain insights into datasets, but the inclusion of additional information about models would enable claims about (model, dataset, task) triples.
Future Work. We envision Rings to support the design of better datasets and more challenging tasks for graph learning, and we hope that the community will use our proposed measures to further scrutinize its datasets. Beyond these aspirations, we highlight three concrete avenues for future work. First, although we provide some initial theoretical results on the behavior of perturbed mode complementarity, a comprehensive theoretical analysis of its properties is yet outstanding. Second, while our work focused on graph-level tasks, node-level and edge-level tasks also merit closer inspection, and these tasks might naturally give rise to a localized notion of mode complementarity. And third, studying how mode complementarity changes during training, or how mode-complementarity distributions impact train-eval-test splits, could yield interesting insights into the relationship between mode complementarity and performance.
@@ A.3.2 Complementarity under Perturbation
Definition A.7 (Complimentarity for Disconnected Graphs).
@@
$\operatorname{\gamma}_{cg}^{1,1}(G,X)$ $=\operatorname{\mathcal{C}}_{1,1}(D^{\operatorname{\operatorname{\varphi}_{cg}}}_{\operatorname{\operatorname{d}_{S}}},D_{\operatorname{\operatorname{d}_{F}}})$
$=\operatorname{\mathcal{C}}_{1,1}(\operatorname{\mathbf{1}}_{n}-\operatorname{\mathbf{I}}_{n},D_{\operatorname{\operatorname{d}_{F}}})$
$=\left\|(\operatorname{\mathbf{1}}_{n}-\operatorname{\mathbf{I}}_{n})-D_{\operatorname{\operatorname{d}_{F}}}\right\|_{1,1}$
$=\frac{1}{n^{2}-n}\left(\sum_{i=1}^{n}\sum_{j\neq i}^{n}\left|1-\operatorname{\operatorname{d}_{F}}(i,j)\right|\right)$
$=\frac{1}{n^{2}-n}\left(n^{2}-n-\sum_{i=1}^{n}\sum_{j\neq i}^{n}\operatorname{\operatorname{d}_{F}}(i,j)\right)$
$=1-\frac{1}{n^{2}-n}\left(\sum_{i=1}^{n}\sum_{j=1}^{n}\operatorname{\operatorname{d}_{F}}(i,j)\right)$
$=1-\left\|D_{\operatorname{\operatorname{d}_{F}}}\right\|_{1,1}$
$=1-\operatorname{\gamma}_{eg}^{1,1}(G,X)\qed$
The justifications for the above steps are as follows:
  • Line 1: Definition 2.13.
  • Line 2: Lemma A.9.
  • Line 3: Definition 2.11.
  • Line 4: Definition A.6.
  • Line 5: Unit-diameter assumption: $\operatorname{\operatorname{d}_{F}}\leq 1$.
  • Line 6: Axiom (ii) of Definition A.1.
  • Line 7: Definition A.6.
  • Line 8: Definition of $\operatorname{\gamma}_{eg}^{1,1}(G,X)$ and Lemma A.8.
Note that a symmetrical argument can be applied to $\operatorname{\gamma}_{cf}^{1,1}$. This concludes the proof.
@@ C.2.1 Molecules and Compounds
The AIDS dataset (Riesen & Bunke, 2008), for example, is built from a repository of molecules in the AIDS Antiviral Screen Database of Active Compounds (NIH National Cancer Institute, 2004). Nodes have four features (Morris et al., 2020), most notably their unique atomic number. This is a binary-classification dataset in which molecules are classified by whether or not they demonstrate activity against the HIV virus. Among the 2 000 graphs in this dataset, note that the two classes are imbalanced: there are four times as many inactive molecules (1 600) as active (400). The MolHIV (ogbg-molhiv) dataset, introduced by the Open Graph Benchmark (Hu et al., 2020) with over 40,000 molecules from MoleculeNet (Wu et al., 2018), proposes a similar task. Like AIDS, MolHIV classifies molecules into two classes based on their ability or inability to suppress the HIV virus. However, nodes in the MolHIV dataset have $9$ features (compared to $4$ in AIDS), including atomic number, chirality, and formal charge.
@@ D.1 Extended Setup
for each $(\operatorname{\Phi}(D),\operatorname{\mathcal{A}})$
@@
Evaluation Strategy Tuned model $(\operatorname{\Phi}(D),\operatorname{\mathcal{A}},\operatorname{\theta})$ re-trained
@@
then evaluated on $\operatorname{\Phi}(D)$’s test set.
Table 4: GNN setup to evaluate performance separability. Parameter combinations are specified in Table 5. $\operatorname{\Phi}_{rg}(\operatorname{\mathcal{D}})$ perturbations were trained using a $\mathcal{R}_{F,1}(\cdot,10)$ randomization. The top performing parameter combination $\operatorname{\theta}^{\star}$ was selected based on validation statistics: primarily based on AUROC, with ties broken by accuracy and then loss.
@@
Table 5: GNN tuning parameters. For consistency, all $(\operatorname{\Phi}(\operatorname{\mathcal{D}}),\operatorname{\mathcal{A}})$ were tuned across a consistent hyperparameter grid search.
@@ D.2.2 Performance-Separability Results
Reasoning first principles, we assert that the most instructive graph-learning datasets should have a separable original mode. In other words, we would like both the features and the graph structure to be informative and necessary to solve a given task. When a perturbed mode either outperforms or is non-separable from the original, then we can say that the perturbed mode is non-essential for the task at hand.

### v2

@@ Abstract
Benchmark datasets have proved pivotal to the success of graph learning, and good benchmark datasets are crucial to guide the development of the field. Recent research has highlighted problems with graph-learning datasets and benchmarking practices—revealing, for example, that methods which ignore the graph structure can outperform graph-based approaches. Such findings raise two questions: (1) What makes a good graph-learning dataset, and (2) how can we evaluate dataset quality in graph learning? Our work addresses these questions. As the classic evaluation setup uses datasets to evaluate models, it does not apply to dataset evaluation. Hence, we start from first principles. Observing that graph-learning datasets uniquely combine two modes—graph structure and node features—, we introduce Rings, a flexible and extensible mode-perturbation framework to assess the quality of graph-learning datasets based on dataset ablations—i.e., quantifying differences between the original dataset and its perturbed representations. Within this framework, we propose two measures—performance separability and mode complementarity—as evaluation tools, each assessing the capacity of a graph dataset to benchmark the power and efficacy of graph-learning methods from a distinct angle. We demonstrate the utility of our framework for dataset evaluation via extensive experiments on graph-level tasks and derive actionable recommendations for improving the evaluation of graph-learning methods. Our work opens new research directions in data-centric graph learning, and it constitutes a step toward the systematic evaluation of evaluations.
@@ 1 Introduction
    Experiments. We demonstrate the value of our framework and measures through extensive experiments, focusing on real-world datasets with graph-level tasks.
@@ 2.1 Mode Perturbations
Given a dataset $\operatorname{\mathcal{D}}$ of attributed graphs and a mode perturbation $\varphi$, a dataset perturbation is given by ${\operatorname{\varphi}(\operatorname{\mathcal{D}}):=\{\operatorname{\varphi}(G,X):(G,X)\in\operatorname{\mathcal{D}}\}}$.
@@ 2.2 Performance Separability
Intuitively, given two perturbations $\operatorname{\varphi}$, $\operatorname{\varphi}^{\prime}$ of a dataset $\operatorname{\mathcal{D}}$ as well as a task to be solved on $\operatorname{\mathcal{D}}$, performance separability measures the distance between performance distributions associated with $\operatorname{\varphi}(\operatorname{\mathcal{D}})$ and $\operatorname{\varphi}^{\prime}(\operatorname{\mathcal{D}})$. For a formal definition, we need notation describing these distributions.
@@
To elucidate the relationship between the graph structure and the node features as it pertains to performance, within Rings, we tune models not only on datasets $\operatorname{\mathcal{D}}$ but also on perturbed datasets $\operatorname{\varphi}(\operatorname{\mathcal{D}})$.
@@
For a mode perturbation $\operatorname{\varphi}$, $\operatorname{\operatorname{\mathcal{M}}_{\operatorname{\varphi}}}\coloneqq(\operatorname{\varphi}(\operatorname{\mathcal{D}}),\operatorname{\mathcal{A}},\operatorname{\theta})$ denotes a model tuned to solve the task associated with $\operatorname{\mathcal{D}}$ under mode perturbation $\operatorname{\varphi}$.
@@ 3 Related Work
Figure 3: Comparing GNN performance across different versions of the same dataset. We show the mean (dot) and $95$th percentile interval (bars) of accuracy and AUROC across $100$ runs of the best (as measured by the respective performance mean) among our tuned GAT, GCN, and GIN models, for the original version and $5$ perturbations of $13$ graph-learning datasets. Black resp. silver horizontal lines show best mean trained resp. untrained performance on the original data. For Reddit-M, the complete-graph and complete-features perturbations failed to train due to memory problems, but the existing results already allow us to conclude that this dataset lacks performance separability. Note that AUROC is the recommended evaluation metric for MolHIV; we include accuracy here only for completeness.
@@ 4 Experiments
We now demonstrate how to use Rings to evaluate the quality of graph-learning datasets, guided by the principles developed in Section 1. Leveraging our mode-perturbation framework, we explore P1 via performance separability and P2 via mode complementarity, before distilling our observations into an actionable taxonomy of recommendations. For further details and additional results, including regression tasks, transformer architectures, and graph-level performance comparisons between models, see Appendix D. [^3] Our reproducibility package is publicly available via Zenodo at 10.5281/zenodo.15547322. The code is maintained on GitHub at https://github.com/aidos-lab/rings.
Evaluated Datasets. In our main experiments, we evaluate $13$ popular graph-classification datasets: From the life sciences, we select AIDS, ogbg-molhiv (MolHIV), MUTAG, and NCI1 (small molecules), as well as DD, Enzymes, Peptides-func (Peptides), and PROTEINS-full (Proteins) (larger chemical structures). From the social sciences, we take COLLAB, IMDB-B, and IMDB-M (collaboration ego-networks), as well as REDDIT-B and REDDIT-M (online interactions). To approximate the standard evaluation setup as closely as possible and ensure compatibility with a wide range of models, for all datasets, we use the node features as encoded by PyTorch-geometric and disregard the edge features. See Appendix C for details and references.
@@ 4.1 Performance Separability (P1)
For each version of each dataset, we identify the model with the best performance under evaluation metric $\operatorname{f}\in$accuracy, AUROC} as $(\operatorname{\mathcal{A}}^{\star},\operatorname{\theta}^{\star})\coloneqq\arg\max_{(\operatorname{\mathcal{A}},\operatorname{\theta})}\mathbb{E}\left[\operatorname{\operatorname{f}(\operatorname{\operatorname{\mathcal{M}}_{\operatorname{\varphi}}};\operatorname{\zeta})}\right]$, using the performance distribution of the top model, $\operatorname{\mathcal{M}}^{\star}_{\operatorname{\varphi}}\coloneqq(\operatorname{\varphi}(\operatorname{\mathcal{D}}),\operatorname{\mathcal{A}}^{\star},\operatorname{\theta}^{\star})$, to assess performance separability.
@@ 4.3 Dataset Taxonomy
Taxonomy, categories, and interpretation. Our taxonomy is built on our measures of performance separability and structural diversity (derived from mode complementarity). Here, we emphasize structural diversity over feature diversity because the structural mode is characteristic of the graph-learning setting, acknowledging that datasets with high feature diversity could still be useful as benchmarks in settings outside of graph learning. Here, we call the value of a measure high if it is judged as $\circ$, $+$, or +++++ + (see Tables 1 and 2), and low otherwise. In the following, we elaborate on the definition and interpretation of each action category.
[…]
Separability, complementarity, and misalignment. While our experiments show that higher mode complementarity is associated with higher performance (see Figure 4), we make no claims about the relationship between mode diversity and performance separability. Since performance separability assesses the task-specific performance gap between a dataset and its perturbations, and mode diversity measures the task-agnostic variation contained in the modes of an individual dataset, we would not expect high mode diversity to be consistently associated with high performance separability. For example, a dataset could have high structural and feature diversity but low mode complementarity, which could lead to complete graph and complete features performing on par with the original mode, eliminating separability.
Misalignment occurs when there is interesting variation in the data that models could leverage in their predictions (high mode diversity), but the existing relationship between this variation and the prediction target does not provide a significant performance advantage over settings in which this relationship is deliberately destroyed (lack of performance separability). While the lack of performance separability could also be due to limitations of the models employed (i.e., they may not be expressive enough), given our comprehensive measurement setup, we deem it more likely that the relationship between the variation in the data and the prediction target is strained, prompting the need for realignment. However, more work is needed to disentangle the model-related and the data-related factors contributing to a lack of performance separability.
@@ 5 Discussion
Conclusions. We introduced Rings, a principled framework for evaluating graph-learning datasets based on mode perturbations—i.e., controlled changes to their graph structure and node features. In Rings, we developed performance separability as a task-dependent, and mode complementarity as a task-independent measure of dataset quality. Using our framework to categorize $13$ popular graph-classification datasets, we identified several benchmarks that require realignment (e.g., by better data modeling) or complete deprecation.
While some datasets have been observed to be problematic in prior work, with Rings, we offer a full perturbation-based evaluation pipeline. We designed our initial set of perturbations from a data-centric perspective, but additional, cleverly crafted perturbations could also isolate and confirm the performance separability of new model contributions. Thus, we hope that Rings will raise the standard of proof for model-centric evaluation in graph learning as well.
@@
Limitations. Given that message passing is the predominant paradigm in graph learning, our performance-separability experiments mostly targeted message-passing GNNs, at the exclusion of other architectures. While Rings is a general framework, our current norm-based comparison of metric spaces assumes that nodes and node features are paired. Thus, we presently do not measure aspects like the metric distortion arising from different mappings. Moreover, our notion of mode complementarity is task-independent and model-agnostic. This allows us to gain insights into datasets, but the inclusion of additional information about models would enable claims about (model, dataset, task) triples.
Future Work. We envision Rings to support the design of better datasets and more challenging tasks for graph learning, and we hope that the community will use our measures to further scrutinize its evaluation infrastructure. Additionally, we highlight four concrete avenues for future work.
  1. F1
    Theoretical analysis. While we provided initial theoretical results on the behavior of perturbed mode complementarity, a comprehensive theoretical analysis of its properties is yet outstanding. Adopting an information-theoretic perspective could yield valuable insights.
  2. F2
    Task coverage. Our work focused on graph-level tasks, but node-level tasks and edge-level tasks, too, merit closer inspection. While performance separability immediately applies to such tasks, extending mode complementarity requires further conceptual advances. These advances might also naturally give rise to a localized notion of mode complementarity.
  3. F3
    Feature coverage and graph coverage. Rings naturally accounts for node features, but some datasets also contain edge features or graph-level features that could be incorporated into our framework. Likewise, defining mode complementarity for more expressive types of graphs—from temporal graphs to hypergraphs—would constitute a valuable extension of Rings.
  4. F4
    Model interpretability.  While Rings was designed for dataset evaluation, it can also enhance the interpretability of graph-learning models. For example, one could study how mode complementarity changes during training, or how mode-complementarity distributions impact train-eval-test splits. This could yield interesting insights into the relationship between mode complementarity and performance, and it could help elucidate the inner workings of individual models.
Finally, Rings could support the design of better graph-learning models. As highlighted by our dataset taxonomy, by identifying datasets that do not require the specific capabilities of graph-learning models, performance separability helps direct model-development resources to where they are most needed. Moreover, mode complementarity and mode diversity could guide architectural choices for a specific task on a specific dataset (e.g., whether a model should focus on leveraging the graph structure or the node features). These measures could also inform model designs that incorporate data-centric components to enhance performance adaptively (both at the level of individual observations and at the level of entire datasets)—e.g., by preprocessing the data to increase mode complementarity. However, more research is needed to realize the potential of Rings in these domains.
Acknowledgments
This work has received funding from the Swiss State Secretariat for Education, Research, and Innovation (SERI). We acknowledge the use of the High Performance Computing cluster at Helmholtz Munich, maintained by the Scientific Computing and DigIT teams, as well as the computational resources provided by the Aalto Science-IT project.
@@ A.3.2 Complementarity under Perturbation
Definition A.7 (Complementarity for Disconnected Graphs).
@@
[…]
=$\overset{\operatorname{\operatorname{d}_{F}}\leq 1}{=}\frac{1}{n^{2}-n}\left(n^{2}-n-\sum_{i=1}^{n}\sum_{j\neq i}^{n}\operatorname{\operatorname{d}_{F}}(i,j)\right)$
[…]
A symmetrical argument can be applied to $\operatorname{\gamma}_{cf}^{1,1}$, concluding the proof.
@@ C.2.1 Molecules and Compounds
The AIDS dataset (Riesen & Bunke, 2008), for example, is built from a repository of molecules in the AIDS Antiviral Screen Database of Active Compounds (NIH National Cancer Institute, 2004). Nodes have four features (Morris et al., 2020), most notably their unique atomic number. This is a binary-classification dataset in which molecules are classified by whether or not they demonstrate activity against the HIV virus. Among the 2 000 graphs in this dataset, note that the two classes are imbalanced: there are four times as many inactive molecules (1 600) as active (400). The MolHIV (ogbg-molhiv) dataset, introduced by the Open Graph Benchmark (Hu et al., 2020) with over 40,000 molecules from MoleculeNet (Wu et al., 2018b), proposes a similar task. Like AIDS, MolHIV classifies molecules into two classes based on their ability or inability to suppress the HIV virus. However, nodes in the MolHIV dataset have $9$ features (compared to $4$ in AIDS), including atomic number, chirality, and formal charge.
@@ D.1 Extended Setup
for each $(\operatorname{\varphi}(D),\operatorname{\mathcal{A}})$
@@
Evaluation Strategy Tuned model $(\operatorname{\varphi}(D),\operatorname{\mathcal{A}},\operatorname{\theta})$ re-trained
@@
then evaluated on $\operatorname{\varphi}(D)$’s test set.
Table 4: GNN setup to evaluate performance separability. Parameter combinations are specified in Table 5. $\operatorname{\varphi}_{rg}(\operatorname{\mathcal{D}})$ perturbations were trained using a $\mathcal{R}_{F,1}(\cdot,10)$ randomization. The top performing parameter combination $\operatorname{\theta}^{\star}$ was selected based on validation statistics: primarily based on AUROC, with ties broken by accuracy and then loss.
@@
Table 5: GNN tuning parameters. For consistency, all $(\operatorname{\varphi}(\operatorname{\mathcal{D}}),\operatorname{\mathcal{A}})$ were tuned across a consistent hyperparameter grid search.
@@ D.2.2 Performance-Separability Results
Reasoning from first principles, we assert that the most instructive graph-learning datasets should have a separable original mode. In other words, we would like both the features and the graph structure to be informative and necessary to solve a given task. When a perturbed mode either outperforms or is non-separable from the original, then we can say that the perturbed mode is non-essential for the task at hand.